# Entanglement and Bell Nonlocality in $\tau^{+}\tau^{-}$ at the LHC using Machine Learning for Neutrino Reconstruction

### v1

@@ 6 Results
We see that both entanglement and Bell nonlocality can be observed with a significance well above $5\sigma$. The precision, accounting for statistical and systematic uncertainties, on the concurrence is $6.5\%$ with the decay method and $<0.1\%$ with the kinematic method. The measurement of Bell nonlocality has a precision of $2.1\%$ with the decay method and $<0.1\%$ with the kinematic method. These precise measurements, already with the current dataset, highlight the $\tau^{+}\tau^{-}$ channel as likely the first two qubit state at the LHC to observe Bell nonlocality.
@@
$B^{-}_{n}$ $-0.01^{+0.00}_{-0.00}$ $\phantom{-}0.01^{+0.00}_{-0.00}$ $\phantom{-}0.00^{+0.00}_{-0.00}$ $-0.01^{+0.00}_{-0.00}$ $\phantom{-}0.01^{+0.00}_{-0.00}$ $\phantom{-}0.00^{+0.00}_{-0.00}$ $\phantom{-}0.00^{+0.00}_{-0.00}$ $\phantom{-}0.00^{+0.00}_{-0.00}$ $\phantom{-}0.00^{+0.00}_{-0.00}$ $\phantom{-}0.00^{+0.00}_{-0.00}$ $\phantom{-}0.00^{+0.00}_{-0.00}$ $\phantom{-}0.00^{+0.00}_{-0.00}$
$B^{-}_{r}$ $\phantom{-}0.00^{+0.00}_{-0.00}$ $-0.01^{+0.00}_{-0.00}$ $-0.00^{+0.00}_{-0.00}$ $\phantom{-}0.00^{+0.00}_{-0.00}$ $-0.01^{+0.00}_{-0.00}$ $-0.00^{+0.00}_{-0.00}$ $-0.00^{+0.00}_{-0.00}$ $-0.00^{+0.00}_{-0.00}$ $-0.00^{+0.00}_{-0.00}$ $-0.00^{+0.00}_{-0.00}$ $-0.00^{+0.00}_{-0.00}$ $-0.00^{+0.00}_{-0.00}$
@@
$B^{+}_{r}$ $\phantom{-}0.00^{+0.00}_{-0.00}$ $\phantom{-}0.01^{+0.00}_{-0.00}$ $\phantom{-}0.00^{+0.00}_{-0.00}$ $\phantom{-}0.00^{+0.00}_{-0.00}$ $\phantom{-}0.01^{+0.00}_{-0.00}$ $\phantom{-}0.00^{+0.00}_{-0.00}$ $-0.00^{+0.00}_{-0.00}$ $-0.00^{+0.00}_{-0.00}$ $-0.00^{+0.00}_{-0.00}$ $-0.00^{+0.00}_{-0.00}$ $-0.00^{+0.00}_{-0.00}$ $-0.00^{+0.00}_{-0.00}$
@@
$C_{kn}$ $-0.01^{+0.00}_{-0.00}$ $\phantom{-}0.03^{+0.00}_{-0.00}$ $\phantom{-}0.00^{+0.00}_{-0.00}$ $-0.01^{+0.00}_{-0.00}$ $\phantom{-}0.03^{+0.00}_{-0.00}$ $-0.00^{+0.00}_{-0.00}$ $\phantom{-}0.00^{+0.00}_{-0.00}$ $\phantom{-}0.00^{+0.00}_{-0.00}$ $\phantom{-}0.00^{+0.00}_{-0.00}$ $\phantom{-}0.00^{+0.00}_{-0.00}$ $\phantom{-}0.00^{+0.00}_{-0.00}$ $\phantom{-}0.00^{+0.00}_{-0.00}$
$C_{kr}$ $\phantom{-}0.01^{+0.00}_{-0.00}$ $-0.00^{+0.00}_{-0.00}$ $-0.00^{+0.00}_{-0.00}$ $\phantom{-}0.01^{+0.00}_{-0.00}$ $-0.00^{+0.00}_{-0.00}$ $-0.01^{+0.00}_{-0.00}$ $\phantom{-}0.00^{+0.00}_{-0.00}$ $\phantom{-}0.00^{+0.00}_{-0.00}$ $\phantom{-}0.00^{+0.00}_{-0.00}$ $\phantom{-}0.01^{+0.00}_{-0.00}$ $\phantom{-}0.01^{+0.00}_{-0.00}$ $\phantom{-}0.01^{+0.00}_{-0.00}$
@@
$C_{nk}$ $\phantom{-}0.00^{+0.00}_{-0.00}$ $-0.00^{+0.00}_{-0.00}$ $-0.00^{+0.00}_{-0.00}$ $\phantom{-}0.00^{+0.01}_{-0.01}$ $-0.00^{+0.00}_{-0.00}$ $-0.00^{+0.00}_{-0.00}$ $\phantom{-}0.00^{+0.00}_{-0.00}$ $\phantom{-}0.00^{+0.00}_{-0.00}$ $\phantom{-}0.00^{+0.00}_{-0.00}$ $\phantom{-}0.00^{+0.00}_{-0.00}$ $\phantom{-}0.00^{+0.00}_{-0.00}$ $\phantom{-}0.00^{+0.00}_{-0.00}$
@@ 6.1 Systematics Impact
Table 8: Summary of systematic uncertainty impacts on the measurement of concurrence and Bell nonlocality. The table presents results for the two most sensitive channels, $\mu\rho$ and $\rho\rho$, as well as the combined result of all 7 subchannels. The uncertainties are evaluated in the signal region (SR Only) and with trigger effects included (SR & Trigger). The total systematic uncertainty is computed as the quadrature sum of individual contributions.
@@
$\mu\rho$ $\rho\rho$ Combined $\mu\rho$ $\rho\rho$ Combined
All Systematics 4.96% 3.62% 12.85% 8.40% 4.29% 42.20%
MC Statistics 4.88% 3.59% 12.35% 7.84% 4.25% 38.57%
Luminosity 0.02% $<0.01$% 0.37% 2.10% 0.11% 8.21%
Background Cross-Section 0.56% $<0.01$% 0.82% 0.41% $<0.01$% 1.43%
Signal Cross-Section 0.04% 0.03% 0.70% 0.85% 0.07% 3.24%
Tau Energy Scale 0.24% 0.30% 0.87% 0.33% 0.13% 1.98%
Jet Enery Scale 0.28% 0.18% 1.84% 0.65% 0.25% 10.84%
Soft MET ($p_{x}$, $p_{y}$) 0.61% 0.23% 2.70% 1.81% 0.49% 9.57%
$\nu$ Sampling $<0.01$% $<0.01$% 0.01% 0.02% $<0.01$% 0.10%
@@ 8 Acknowledgements
The authors would like to thank Kun Cheng, Navin McGinnis, and Vinicius Mikuni for helpful conversations. YZ and SCH are supported by the U.S. National Science Foundation under Grant Number 2209034. Zhou and Li are supported by National Key R&D Program of China (Nos. 2023YFA1605703 and 2023YFA1605700), Shanghai Pilot Program for Basic Research – Shanghai Jiao Tong University (No. 21TQ1400209), National Natural Science Foundation of China (No. 12150006). Zhou is supported by T.D. Lee scholarship. This work was supported in part by the US Department of Energy under grant No. DE-SC0007914 and in part by Pitt PACC. ML is also supported by the National Science Foundation under grant No. PHY-2112829.
@@ Appendix A Full Density Matrix
with $\mathbb{R}^{q\bar{q}}(\hat{s},\theta)=R^{q\bar{q}}(\hat{s},\theta)+R^{q\bar{q}}(\hat{s},\pi-\theta),$ (23)
@@ B.3 Density Matrix
$B^{-}_{n}$ $\phantom{-}0.00^{+0.00}_{-0.00}$ $-0.00^{+0.02}_{-0.02}$ $-0.01^{+0.06}_{-0.06}$ $-0.00^{+0.14}_{-0.13}$ $\phantom{-}0.00^{+0.31}_{-0.31}$ $-0.00^{+0.06}_{-0.06}$ $\phantom{-}0.01^{+0.02}_{-0.02}$ $\phantom{-}0.00^{+0.01}_{-0.01}$ $-0.00^{+0.04}_{-0.04}$ $-0.01^{+0.06}_{-0.06}$ $-0.00^{+0.11}_{-0.11}$ $\phantom{-}0.00^{+0.30}_{-0.30}$ $-0.00^{+0.06}_{-0.06}$ $\phantom{-}0.01^{+0.02}_{-0.02}$
$B^{-}_{r}$ $\phantom{-}0.00^{+0.01}_{-0.01}$ $\phantom{-}0.00^{+0.01}_{-0.01}$ $\phantom{-}0.00^{+0.02}_{-0.02}$ $-0.00^{+0.05}_{-0.05}$ $-0.01^{+0.07}_{-0.07}$ $\phantom{-}0.00^{+0.09}_{-0.09}$ $-0.01^{+0.02}_{-0.02}$ $\phantom{-}0.00^{+0.04}_{-0.03}$ $\phantom{-}0.00^{+0.02}_{-0.02}$ $\phantom{-}0.00^{+0.03}_{-0.03}$ $-0.00^{+0.04}_{-0.04}$ $-0.01^{+0.07}_{-0.07}$ $\phantom{-}0.00^{+0.10}_{-0.10}$ $-0.01^{+0.02}_{-0.02}$
@@
$B^{+}_{n}$ $\phantom{-}0.00^{+0.00}_{-0.00}$ $-0.00^{+0.24}_{-0.23}$ $\phantom{-}0.00^{+0.01}_{-0.01}$ $\phantom{-}0.00^{+0.02}_{-0.02}$ $\phantom{-}0.00^{+0.02}_{-0.02}$ $\phantom{-}0.00^{+0.02}_{-0.02}$ $\phantom{-}0.00^{+0.02}_{-0.02}$ $\phantom{-}0.00^{+0.02}_{-0.01}$ $-0.00^{+0.70}_{-0.44}$ $\phantom{-}0.00^{+0.02}_{-0.02}$ $\phantom{-}0.00^{+0.03}_{-0.03}$ $\phantom{-}0.00^{+0.02}_{-0.02}$ $\phantom{-}0.00^{+0.03}_{-0.03}$ $\phantom{-}0.00^{+0.02}_{-0.02}$
$B^{+}_{r}$ $\phantom{-}0.00^{+0.01}_{-0.01}$ $-0.00^{+0.02}_{-0.02}$ $\phantom{-}0.00^{+0.00}_{-0.00}$ $\phantom{-}0.00^{+0.01}_{-0.01}$ $\phantom{-}0.00^{+0.01}_{-0.01}$ $\phantom{-}0.01^{+0.00}_{-0.00}$ $\phantom{-}0.01^{+0.00}_{-0.00}$ $\phantom{-}0.00^{+0.09}_{-0.03}$ $-0.00^{+0.04}_{-0.03}$ $\phantom{-}0.00^{+0.00}_{-0.00}$ $\phantom{-}0.00^{+0.01}_{-0.01}$ $\phantom{-}0.00^{+0.01}_{-0.01}$ $\phantom{-}0.01^{+0.00}_{-0.00}$ $\phantom{-}0.01^{+0.00}_{-0.00}$
@@
$C_{kn}$ $\phantom{-}0.00^{+0.01}_{-0.01}$ $\phantom{-}0.00^{+0.06}_{-0.06}$ $-0.01^{+0.10}_{-0.10}$ $-0.00^{+0.02}_{-0.02}$ $\phantom{-}0.01^{+0.09}_{-0.09}$ $\phantom{-}0.02^{+0.30}_{-0.31}$ $\phantom{-}0.03^{+0.34}_{-0.35}$ $\phantom{-}0.00^{+0.16}_{-0.13}$ $\phantom{-}0.00^{+0.29}_{-0.25}$ $-0.01^{+0.16}_{-0.16}$ $-0.00^{+0.02}_{-0.02}$ $\phantom{-}0.01^{+0.06}_{-0.06}$ $\phantom{-}0.02^{+0.20}_{-0.21}$ $\phantom{-}0.03^{+0.29}_{-0.29}$
$C_{kr}$ $-0.00^{+0.00}_{-0.00}$ $-0.00^{+0.06}_{-0.06}$ $\phantom{-}0.01^{+0.12}_{-0.12}$ $-0.00^{+0.14}_{-0.14}$ $-0.01^{+0.01}_{-0.01}$ $\phantom{-}0.01^{+0.19}_{-0.19}$ $-0.00^{+0.26}_{-0.26}$ $-0.00^{+0.06}_{-0.04}$ $-0.00^{+0.35}_{-0.29}$ $\phantom{-}0.01^{+0.17}_{-0.17}$ $-0.00^{+0.10}_{-0.10}$ $-0.01^{+0.01}_{-0.01}$ $\phantom{-}0.01^{+0.15}_{-0.15}$ $-0.00^{+0.19}_{-0.19}$
@@
$C_{rk}$ $-0.00^{+0.01}_{-0.01}$ $\phantom{-}0.00^{+0.09}_{-0.09}$ $\phantom{-}0.01^{+0.09}_{-0.09}$ $-0.00^{+0.06}_{-0.06}$ $-0.00^{+0.05}_{-0.05}$ $-0.03^{+0.24}_{-0.24}$ $\phantom{-}0.00^{+0.01}_{-0.01}$ $-0.00^{+0.17}_{-0.06}$ $\phantom{-}0.00^{+0.19}_{-0.16}$ $\phantom{-}0.01^{+0.21}_{-0.20}$ $-0.00^{+0.08}_{-0.08}$ $-0.00^{+0.05}_{-0.05}$ $-0.03^{+0.43}_{-0.44}$ $\phantom{-}0.00^{+0.01}_{-0.01}$
@@
$C_{nk}$ $-0.00^{+0.03}_{-0.03}$ $\phantom{-}0.00^{+0.09}_{-0.09}$ $\phantom{-}0.00^{+0.24}_{-0.24}$ $\phantom{-}0.00^{+0.05}_{-0.05}$ $\phantom{-}0.00^{+0.11}_{-0.11}$ $-0.01^{+0.43}_{-0.44}$ $-0.00^{+0.01}_{-0.01}$ $-0.00^{+0.52}_{-0.16}$ $\phantom{-}0.00^{+0.34}_{-0.20}$ $\phantom{-}0.00^{+0.68}_{-0.65}$ $\phantom{-}0.00^{+0.06}_{-0.06}$ $\phantom{-}0.00^{+0.13}_{-0.12}$ $-0.01^{+0.56}_{-0.58}$ $-0.00^{+0.01}_{-0.01}$

### v2

@@ 6 Results
We see that both entanglement and Bell nonlocality can be observed with a significance well above $5\sigma$. The precision, accounting for statistical and systematic uncertainties, on the concurrence is $6.5\%$ with the decay method and $0.1\%$ with the kinematic method. The measurement of Bell nonlocality has a precision of $2.1\%$ with the decay method and $0.1\%$ with the kinematic method. These precise measurements, already with the current dataset, highlight the $\tau^{+}\tau^{-}$ channel as likely the first two qubit state at the LHC to observe Bell nonlocality.
@@
[…]
@@
$B^{+}_{r}$ $\phantom{-}0.00^{+0.00}_{-0.00}$ $\phantom{-}0.01^{+0.00}_{-0.00}$ $\phantom{-}0.00^{+0.00}_{-0.00}$ $\phantom{-}0.00^{+0.00}_{-0.00}$ $\phantom{-}0.01^{+0.00}_{-0.00}$ $\phantom{-}0.00^{+0.00}_{-0.00}$ $^{+0.00}_{-0.00} 0.000.00$$^{+0.00}_{-0.00}$ $^{+0.00}_{-0.00} 0.000.00$$^{+0.00}_{-0.00}$ $^{+0.00}_{-0.00} 0.000.00$$^{+0.00}_{-0.00}$
@@
[…]
@@
$C_{nk}$ $\phantom{-}0.00^{+0.00}_{-0.00}$ $^{+0.00}_{-0.00} 0.000.00$$^{+0.00}_{-0.00}$ $\phantom{-}0.00^{+0.01}_{-0.01}$ $^{+0.00}_{-0.00} 0.000.00$$^{+0.00}_{-0.00}$ $\phantom{-}0.00^{+0.00}_{-0.00}$ $\phantom{-}0.00^{+0.00}_{-0.00}$ $\phantom{-}0.00^{+0.00}_{-0.00}$ $\phantom{-}0.00^{+0.00}_{-0.00}$ $\phantom{-}0.00^{+0.00}_{-0.00}$ $\phantom{-}0.00^{+0.00}_{-0.00}$
@@ 6.1 Systematics Impact
Table 8: Summary of systematic uncertainty impacts on the measurement of concurrence and Bell nonlocality, ranked by their relative effects to the total uncertainty. The table presents results for the two most sensitive channels, $\mu\rho$ and $\rho\rho$, as well as the combined result of all 7 subchannels. The uncertainties are evaluated in the signal region (SR Only) and with trigger effects included (SR & Trigger). The total systematic uncertainty is computed as the quadrature sum of individual contributions.
@@
$\rho\rho$ $\mu\rho$ Combined $\rho\rho$ $\mu\rho$ Combined
All Systematics 29.81% 29.76% 31.29% 31.00% 29.82% 31.35%
MC Statistics 29.31% 29.56% 30.05% 28.93% 29.55% 28.66%
Luminosity 0.12% 0.08% 0.90% 7.73% 0.74% 6.10%
Background Cross-Section 3.39% 0.07% 2.01% 1.51% 0.05% 1.06%
Signal Cross-Section 0.23% 0.23% 1.71% 3.12% 0.52% 2.41%
Tau Energy Scale 1.47% 2.50% 2.12% 1.20% 0.89% 1.47%
Jet Enery Scale 1.67% 1.50% 4.49% 2.41% 1.72% 8.05%
Soft MET ($p_{x}$, $p_{y}$) 3.66% 1.90% 6.57% 6.68% 3.42% 7.11%
$\nu$ Sampling 0.02% 0.02% 0.03% 0.06% 0.03% 0.07%
@@ 8 Acknowledgements
The authors would like to thank Quentin Baut, Kun Cheng, Lawrence Lee, Navin McGinnis, and Vinicius Mikuni for helpful conversations. YZ and SCH are supported by the U.S. National Science Foundation under Grant Number 2209034. SCH is partially supported by DGE-2021540. Zhou and Li are supported by National Key R&D Program of China (Nos. 2023YFA1605703 and 2023YFA1605700), Shanghai Pilot Program for Basic Research – Shanghai Jiao Tong University (No. 21TQ1400209), National Natural Science Foundation of China (No. 12150006). Zhou is supported by T.D. Lee scholarship. This work was supported in part by the US Department of Energy under grant No. DE-SC0007914 and in part by Pitt PACC. ML is also supported by the National Science Foundation under grant No. PHY-2112829.
@@ Appendix A Full Density Matrix
with $\hskip-1.99997pt\mathbb{R}^{q\bar{q}}(\hat{s},\theta)=R^{q\bar{q}}(\hat{s},\theta)+R^{q\bar{q}}(\hat{s},\pi-\theta),$ (23)
@@
$\tilde{A}^{q\bar{q}}=$
$\frac{1}{16}\bigg\{Q_{q}^{2}Q_{\tau}^{2}\big[2-\beta^{2}\sin^{2}\theta\big]+2Q_{q}Q_{\tau}\operatorname{Re}\big[\chi(\hat{s})\big]\Big[2\beta g_{A}^{q}g_{A}^{\tau}\cos\theta+g_{V}^{q}g_{V}^{\tau}\big(2-\beta^{2}\sin^{2}\theta\big)\Big]$
$\big|\chi(\hat{s})\big|^{2}\bigg[\big(g_{V}^{q2}+g_{A}^{q2}\big)\Big(2g_{V}^{\tau 2}+2\beta^{2}g_{A}^{\tau 2}-\beta^{2}\big(g_{V}^{\tau 2}+g_{A}^{\tau 2}\big)\sin^{2}\theta\Big)+8\beta g_{V}^{q}g_{V}^{\tau}g_{A}^{q}g_{A}^{\tau}\cos\theta\bigg]\bigg\},$
$\tilde{B}_{k}^{\pm,q\bar{q}}=$
$\frac{1}{8}\Big\{Q_{q}Q_{\tau}\operatorname{Re}\big[\chi(\hat{s})\big]\Big[\beta g_{A}^{\tau}g_{V}^{q}\big(1+\cos^{2}\theta\big)+2g_{A}^{q}g_{V}^{\tau}\cos\theta\Big]$
$\big|\chi(\hat{s})\big|^{2}\Big[2g_{A}^{q}g_{V}^{q}\big(\beta^{2}g_{A}^{\tau 2}+g_{V}^{\tau 2}\big)\cos\theta+\beta g_{A}^{\tau}g_{V}^{\tau}\big(g_{V}^{q2}+g_{A}^{q2}\big)\big(1+\cos^{2}\theta\big)\Big]\Big\},$
$\tilde{B}_{r}^{\pm,q\bar{q}}=$
$\frac{1}{8}\sin\theta\sqrt{1-\beta^{2}}\Big\{Q_{q}Q_{\tau}\operatorname{Re}\big[\chi(\hat{s})\big]\big[\beta g_{A}^{\tau}g_{V}^{q}\cos\theta+2g_{A}^{q}g_{V}^{\tau}\big]$
$\big|\chi(\hat{s})\big|^{2}g_{V}^{\tau}\Big[\beta g_{A}^{\tau}\big(g_{V}^{q2}+g_{A}^{q2}\big)\cos\theta+2g_{A}^{q}g_{V}^{q}g_{V}^{\tau}\Big]\Big\},$
$\tilde{B}_{n}^{\pm,q\bar{q}}=$
 0,\,0,
$\tilde{C}_{nn}^{q\bar{q}}=$
$\frac{1}{16}\beta^{2}\sin^{2}\theta\Big\{Q_{q}^{2}Q_{\tau}^{2}+2Q_{q}Q_{\tau}\operatorname{Re}\big[\chi(\hat{s})\big]g_{V}^{q}g_{V}^{\tau}-\big|\chi(\hat{s})\big|^{2}\big(g_{V}^{q2}+g_{A}^{q2}\big)\big(g_{A}^{\tau 2}-g_{V}^{\tau 2}\big)\Big\},$
$\tilde{C}_{rr}^{q\bar{q}}=$
$\frac{1}{16}\sin^{2}\theta\Big\{\big(\beta^{2}-2\big)Q_{q}^{2}Q_{\tau}^{2}+2Q_{q}Q_{\tau}\operatorname{Re}\big[\chi(\hat{s})\big]g_{V}^{q}g_{V}^{\tau}\big(\beta^{2}-2\big)$
$\big|\chi(\hat{s})\big|^{2}\Big[\beta^{2}\big(g_{A}^{\tau 2}+g_{V}^{\tau 2}\big)-2g_{V}^{\tau 2}\Big]\big(g_{V}^{q2}+g_{A}^{q2}\big)\Big\},$
$\tilde{C}_{kk}^{q\bar{q}}=$
$\frac{1}{16}\bigg\{Q_{q}^{2}Q_{\tau}^{2}\Big[\big(\beta^{2}-2\big)\sin^{2}\theta+2\Big]$
$2Q_{q}Q_{\tau}\operatorname{Re}\big[\chi(\hat{s})\big]\Big[2\beta g_{A}^{q}g_{A}^{\tau}\cos\theta+g_{V}^{q}g_{V}^{\tau}\big((\beta^{2}-2)\sin^{2}\theta+2\big)\Big]$
$\big|\chi(\hat{s})\big|^{2}\bigg[8\beta g_{A}^{q}g_{A}^{\tau}g_{V}^{q}g_{V}^{\tau}\cos\theta+\big(g_{V}^{q2}+g_{A}^{q2}\big)\Big(2g_{V}^{\tau 2}\cos^{2}\theta-\beta^{2}\big(g_{A}^{\tau 2}-g_{V}^{\tau 2}\big)\sin^{2}\theta+2\beta^{2}g_{A}^{\tau 2}\Big)\!\bigg]\!\bigg\},$
$\tilde{C}_{kr}^{q\bar{q}}=$
$\tilde{C}_{rk}^{q\bar{q}}=\frac{1}{8}\sin\theta\sqrt{1-\beta^{2}}\Big\{Q_{q}^{2}Q_{\tau}^{2}\cos\theta+Q_{q}Q_{\tau}\operatorname{Re}\big[\chi(\hat{s})\big]\left[\beta g_{A}^{q}g_{A}^{\tau}+2g_{V}^{q}g_{V}^{\tau}\cos\theta\right]$
$\hskip 30.00005pt+\big|\chi(\hat{s})\big|^{2}\Big[2\beta g_{A}^{q}g_{A}^{\tau}g_{V}^{q}g_{V}^{\tau}+g_{V}^{\tau 2}\big(g_{V}^{q2}+g_{A}^{q2}\big)\cos\theta\Big]\Big\},$
$\tilde{C}_{nr}^{q\bar{q}}=$
$\tilde{C}_{rn}^{q\bar{q}}=\tilde{C}_{nk}^{q\bar{q}}=\tilde{C}_{kn}^{q\bar{q}}=0,$
@@ B.3 Density Matrix
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]